Your boss believes the company's power plant is producing too much air pollution on a typical island. Your boss gives you three choices for dealing with this problem because he/she does not want to deal with it:

  1. You can pay a pollution tax (Carbon Offsets) one time of $13,000,000 immediately.  
  2. You can close the plant and install a power cable from the mainland to the Island. That will cost you $1,000,000 at the end of this year, $3,000,000 at the end of next year and then $750,000 forever for maintenance.  
  3. You can retrofit the plant with scrubbers to reduce the emissions to make the plant green. That will cost $7.5m at the end of this year and $100,000 for 50-years for maintenance.  

Assume that the cost of generating power on the mainland is approximately the same as the cost of generating power at the Island's plant. Assume, this comes as a surprise to you and you, have not saved any money in reserves, and you need to raise capital. Additional information is that market has a 12 percent market risk premium on the power plant with the risk-free rate being 5 percent with a company tax rate of 35 percent.  

Current total raised capital at the power plant:  (This will help you calculate the WACC)

  • Debt – 7,000 outstanding bonds, at 7.5% coupon and 20 years to maturity. These bonds pay interest semiannually and quoted a price of 108 percent of par.  
  • Common Stock -180,000 shares outstanding, selling for $50 per share: Beta .90.  
  • Preferred Stock – 8,000 shares of 5.5 percent preferred stock outstanding, currently selling for $95.00 per share.

Please answer in essay format and provide your Excel document showing all your calculation in appendixes choose the best option for Island. Support your answer with your calculations. Also, to calculations use specified resources, other appropriate scholarly resources, including older articles.
Answer #1
CALCULATION OF WEIGHTED AVERAGE COST OF CAPITAL
Assuming Face value per Bond $1,000
Pv Market Value per bond $1,080 (1000*1.08)
Pmt Semi annual interest $37.50 (1000*(0.075/2)
Nper Number of semi annual period 40 (20*2)
Fv Payment at maturity $1,000
RATE Semi annual yield 3.38% (using RATE function of excelwith Nper=40,Pmt=37.50,Pv=-1080,Fv=1000)
A Annual yield 6.76% (3.38*2)
B Tax Rate 0.35
C=A*(1-B) Cost of debt 4.40%
Cost of Equity =Rf+Beta*(Rm-Rf)
Rf=Risk free rate=5%
Beta=0.90
Rm-Rf=Market RiskPremium=12%
Cost of Equity=5+0.90*12= 15.80%
D Cost of Common Shares 15.80%
E Cost of preferred share 5.79% (5.5/95)*100 5.789474
E Market Value of Debt $7,560,000 (7000*1080)
F Market value of Common shares $9,000,000 (180000*50)
G Market value of preferred shares $760,000 (8000*95)
H=E+F+G Total capital $17,320,000
Wd=E/H Weight of Debt                 0.44
We=F/H Weight of Common Shares                 0.52
Wp=G/H Weight of preferred shares                 0.04
Weighted Average Cost of capital (WACC)=C*Wd+D*We+G*Wp
Weighted Average Cost of capital (WACC) 10.38%
PRESENT VALUE OF THREE ALTERNATIVES:
ALTERNATIVE 1
Pay Pollution Tax
Present Value of Cost of Alternative 1 $13,000,000

CLOSE PLANT AND INSTALL CABLE Present value of 1000000 at the end of this year Present value of 3000000 at the end of next year Value of 750000at the end of every year in perpetuity(at end of year2) Present value of 750000 per year in perpetuity Present Value of Cost of $905,961 (1000000/1.1038) $2,462,297 (3000000/(1.103842)) $7,225,434 (750000/0.1038) $5,930,388 (7225434/(1.1038A2) $9,298,647 L-l+J+K Alternative 2 ALTERNATIVE 3 RETROFIT WITH SCRUBBERS Present Value of $7.5 millionat the end of this year Value of $100000for 50 years Present Value of $100000 for 50years Present Value of Cost of $6,794,709 (7.5million/1.1038) $956,484 (using PV function of excel with Rate-10.38%, Npep50, Pmt-100000) $866,538 (956484/1.1038) P M+N Alternative 3 $7,661,247M+N Alternative 3 $7,661,247 Present Value of Cost $13,000,000 $9,298,647 $7,661,247 ALTERNATIVE 1 ALTERNATIVE 2 ALTERNATIVE 3 Alternative of Retrofit with Scrubbers should be selectec